# How Internal Structure Shapes the Metallicity of Giant Exoplanets

### v1

@@ 4.2 Connection with formation
Extent of the dilute core.
For some high-metallicity cases ($Z\gtrsim 0.5$; e.g., TOI-4010d and WASP-196b), no solutions were found for compact gradients with $m_{\rm dilute}=0.25$ or $0.5$: the interior could not accommodate the required heavy-element fraction unless the gradient was shifted outward to larger $m_{\rm dilute}$. This rules out highly compact structures and implies a dilute core that extends farther in mass and radius.
In more massive planets, a wide non-adiabatic dilute core could substantially increase the heat retained in the deep interior, altering the radius evolution. Constraining $m_{\rm dilute}$ is therefore crucial for massive planets. A well-confined dilute core may have little impact on inflation (the thermal profile remains close to adiabatic), whereas large $m_{\rm dilute}$ values can stabilize significantly more heat. In non-adiabatic cases, the steepness of the gradient may become important too.
@@
Our results, obtained across different structural configurations, remain consistent with previous studies (Thorngren et al., 2016; Teske et al., 2019; Howard et al., 2025b; Chachan et al., 2025). We recover a clear positive correlation between the total heavy-element mass ($M_{Z}$) and planetary mass ($M$), and a weaker yet significant negative correlation between metallicity ($Z$) and $M$, also when normalised by ($Z_{\star}$).
The overall ZZMM trend indicates a gradual decrease in metallicity with increasing mass; however, the behaviour departs from a sharp decrease of estimated metallicities at low masses (betwenn 0.1 and 0.2 MJ). In the mass range of $\sim$0.2–0.6 $M_{\mathrm{J}}$, we observe a drop in metallicity, consistent with the curvature captured by the fitted power law, but even more evident when expressed as an inverse linear relation, as proposed by Chachan et al. (2025). In this regime, the data exhibit substantial scatter and a near-plateau where further increases in mass do not lead to a corresponding decline in metal content, again reminiscent of the behaviour reported by Chachan et al. (2025) above $\sim$0.28 $M_{\mathrm{J}}$. This is also explaining why the rank correlation test expressed lower correlation coefficients and higher p-values. At the high-mass end, we find hints of renewed enrichment, opposite to the expectations from classical runaway gas accretion, which would predict a dilution of heavy elements as the envelope grows.
@@
This apparent enrichment, together with the significant dispersion across the mass–metallicity plane, likely reflects the diversity of formation pathways and interior structures among giant planets. The presence of high-$Z$, low-$M$ objects (potentially with compositions more akin to Neptune-like planets) suggests that distinct formation channels may operate in this transition regime. On the other side, more massive planets (M ¿ 4 $M_{\mathrm{J}}$) appear to be metal-rich. Even if discerning any conclusion from two planets is premature to say the least, this may suggest a secondary accretion of heavy elements through late planetesimal capture or giant impacts after disk dispersal, leading to enhanced metal enrichment in their deep envelopes (Mordasini et al., 2016; Helled & Stevenson, 2017; Venturini & Helled, 2020)
@@
To assess the robustness of the inferred anti-correlation, we examined how the correlation significance changes when progressively removing planets from the mass extremes. We quantified this sensitivity by recalculating the Kendall $\tau$ correlation and corresponding $p$-values after each removal (Fig. 8). Excluding the lowest-mass planets rapidly diminishes the correlation strength for the CE and DCA models (after only two removals), and after six removals only the FM model retains marginal significance. Conversely, removing the most massive planets has a far smaller impact. This behaviour indicates that the observed ZZMM anti-correlation is predominantly shaped by the high-$Z$, low-$M$ population.
These findings also raise a classification question. Planets with $Z\gtrsim 0.5$ and low masses likely reside near the transition between Neptunian and Jovian regimes, where electron degeneracy pressure no longer dominates the MMRR relation (Müller et al., 2024, and references therein). Treating these as a distinct subpopulation may therefore be warranted. Although the current sample size precludes a robust identification of potential breakpoints via multi-linear or orthogonal-distance regressions, as performed by Müller et al. (2024), future datasets will allow this.
@@ 5 Conclusions
  • We confirm that the inferred bulk metallicity for Fully-Mixed (FM) structures is larger than for Core+Envelope (CE) structures, due to the opacity effect that prevails on the density redistribution effect (Guillot, 2005; Müller et al., 2020; Howard et al., 2025b). This, in turn, leads to Dilute Core (DC) structures that yield similar bulk metallicities than CE, independently of the extent of the composition gradient.
  • When increasing the atmospheric metallicity to the dilute core structure by one (DCA) or three times stellar (DCA3), the inferred bulk metallicity gets closer to the one yielded by FM. This shows how atmospheric composition measurements are crucial to help constrain bulk composition.
  • When considering non-adiabatic temperature gradients in DC structures, we find that the inferred bulk metallicity can be increased by up to 35% for a used value of $R_{\rho}=0.4$. However, future studies should explore a larger parameter space and assess the long-term stability of composition gradients.
  • Using a sample of 44 giant exoplanets, we confirm that heavy element mass increases with planetary mass. The slope of the $M_{Z}-M$ relationship does not change with the assumed structure, but only the intercept. This confirms that giant planets continue to accrete heavy elements as they grow.
@@
Future studies that account for additional physical processes (such as convective mixing), and also observe atmospheric compositions will improve characterisation of planetary interiors. Forthcoming measurements from PLATO and Ariel (expected to be launched in 2026 and 2029, respectively) will sharpen bulk metallicity estimates by providing tighter constraints on mass, radius, age, and atmospheric composition. This will allow a better understanding of the interior of giant planets and of their formation pathways.
@@ Acknowledgements.
We thank Stefano Wirth, Simon Müller, Henrik Knierim, and Shay Zucker for the insightful discussions.
@@
Figure 8:
In order to assess the strength and accuracy of the retrieved $Z$ for the sample used, we took advantage of the fact that the used dataset is the same as the one used by Howard et al. (2025b). We thus took the initial sample of 21 exoplanets with the most precise data on Radius ($R_{err}$¡3%), and compared it with the same planets processed with the method explained in Howard et al. (2025b). Results are shown in Fig. 9, where a visual comparison between the trends for the $ZZ_{\star}$ vs $M$ is being shown. To note how the best fit lines vary from the ones retrieved with the full dataset of 44 exoplanets, showing again the variability and unreliability of such curves for model prediction. Howard also discussed in their paper the shift of the slope $\alpha$ between the dataset considering only $R_{err}$¡3%, or for the broader range considered of $R_{err}$¡5%. In the bottom panels of Fig. 9, it is shown how we assessed the difference between the data from this paper compared to the ones from Howard et al. (2025b). While a simple subtraction could provide information about the difference between the two mean values, we also wanted to investigate whether each ranges of values (so accounting for data uncertainty) lie within each other. To do so, a normalisation using the standard deviations of the data has been computed. The normalised difference can be expressed with the following formula:
$\sigma=\frac{\text{Howard}-\text{CE/FM}}{\sqrt{\sigma_{\text{Howard}}^{2}+\sigma_{\text{CE/FM}}^{2}}}.$ (4)
Results regarding this comparison where considered positive. While there is a slight change in the predictive trends between the two structures CE and FM (slight shift in $\beta$ for CE, and in alphaalpha for FM compared to the ones found in Howard’s. The results for the normalised difference ultimately confirmed the similarities between the two retrieved results, and their respective methods; All values lie between a standard deviation of 1 and -1 (or $\sigma$—¡ 1), showing statistical consistency for all the 21 cases compared. Furthermore, the normalised difference does not show any systematic error related to the mass (since there is no trend in the normalised difference $\sigma$). We thus conclude by saying that there is no discouraging difference between the two results (as the difference between the two approaches can be explained to be within the data error range and thus not due to randomness). We then accepted the validity of our approach explained in Section 2.3, and proceeded with the inspection of the additional structures.
Figure 9:
@@ Appendix C Model assumption-driven changes
Appendix D Weighted least-squares (WLS) $\chi^{2}$ test
To assess the reliability of the power-law fits, we performed a weighted least-squares (WLS) $\chi^{2}$ test, which accounts for individual data uncertainties rather than relying solely on $R^{2}$. In WLS, each residual is scaled by its measurement error:

### v2

@@ 4.2 Connection with formation
Extent of the DC.
For some high-metallicity cases ($Z\gtrsim 0.5$; e.g., TOI-4010d and WASP-196b), no solutions were found for compact gradients with $m_{\rm dilute}=0.25$ or 0.5: the interior could not accommodate the required heavy-element fraction unless the gradient was shifted outwards to higher $m_{\rm dilute}$. Although this behaviour might suggest a more extended dilute region in these planets, this interpretation is not uniquely supported by the data. The inferred extent of dilution is strongly model-dependent and can arise simply from our assumption of a fixed $Z_{\rm atm}$, which permits additional heavy elements to be stored in the envelope without altering the internal composition gradient. In more massive planets, a wide non-adiabatic DC could substantially increase the heat retained in the deep interior, altering the radius evolution. Constraining $m_{\rm dilute}$ is therefore crucial for massive planets. A well-confined DC may have little impact on inflation (the thermal profile remains close to adiabatic), whereas high $m_{\rm dilute}$ values can stabilise significantly more heat. In non-adiabatic cases, the steepness of the gradient may also become important.
@@
Our results, obtained across different structural configurations, are consistent with previous studies (Thorngren et al., 2016; Teske et al., 2019; Howard et al., 2025b; Chachan et al., 2025). We recover a clear positive correlation between the total heavy-element mass ($M_{Z}$) and planetary mass ($M$), and a weaker, yet significant, negative correlation between metallicity ($Z$) and $M$, also when normalised by $Z_{\star}$.
@@
The overall ZZMM trend indicates a gradual decrease in planetary metallicity with increasing mass. There is a sharp decrease in $Z$ for low planetary masses (0.2-0.3 $M_{J}$), followed by a flatter behaviour for higher masses. Other studies on mass-metallicity relations have found similar results (Chachan et al., 2025). At the high-mass end, we find hints of renewed enrichment. This enrichment, together with the significant dispersion across the mass–metallicity relation, likely reflects the diversity of formation pathways and interior structures of giant planets. The presence of high-$Z$, low-$M$ objects (potentially with compositions more akin to Neptune-like planets) is consistent with planet formation theory (Helled, 2023). Interestingly, we find that more massive planets (M ¿ 4 $M_{J}$) are also metal-rich. Even if discerning any conclusion from two planets is premature, this may suggest a secondary accretion of heavy elements through late planetesimal capture or giant impacts after disk dispersal, leading to enhanced metal enrichment in their deep envelopes (Mordasini et al., 2016; Helled and Stevenson, 2017; Venturini and Helled, 2020; Shibata et al., 2022).
@@
To assess the robustness of the inferred anti-correlation, we examined how the correlation significance changes when progressively removing planets from the mass extremes. We quantified this sensitivity by recalculating the Kendall $\tau$ correlation and the corresponding $p$-values after each removal for both the $M_{Z}/M$ and $Z/M$ relations (Fig. 7).
We find that removing low-mass planets has a small effect on the inferred $M_{Z}/M$ relation (top row in Fig. 7): the $p$-values remain nearly unchanged for all the models. In contrast, removing high-mass objects systematically increases the $p$-values. These constitute the ’anchor points’ that most strongly support the monotonic trend found with the correlation test. Even after the removal of six planets from either end, the correlation remains statistically significant, confirming that the observed trend is robust.
For the $Z/M$ relation, the behaviour is different. Because $Z$ already scales inversely with mass ($Z=M_{Z}/M$), plotting $Z$ against $M$ effectively introduces an additional mass dependence in the denominator, which leads to a higher inferred $p$-value (Kronmal, 1993; Brett, 2004). The removal of two low-mass planets is sufficient for the CE and DCA models to lose significance at the $p=0.05$ level. After the removal of six planets, only the FM model retains a marginally significant $p$-value. In contrast, removing the most massive planets has a rather small impact. We find that removing high-mass planets from the sample increases the $p$-values for $M_{Z}$ versus $M$, while removing low-mass planets increases the $p$-values for $Z$ versus $M$. This is because $Z=M_{Z}/M$.
Our findings raise a classification question. Planets with $Z\gtrsim 0.5$ and low masses likely reside near the transition between Neptunian and Jovian regimes, where electron degeneracy pressure no longer dominates the MMRR relation (Müller et al., 2024). The contrast in how the $p$-values increase when removing either low-$M$ or high-$M$ planets could hold a physical meaning. This may indicate the diversity of the sample and the potential existence of sub-populations that do not follow a monotonic trend. In this context, the assumption of a universal, continuous mass–metallicity trend is clearly oversimplified and cannot reflect the large diversity in planetary composition.
@@ 5 Conclusions
  • We confirm that the inferred bulk metallicity for FM structures is higher than for CE structures, due to the opacity effect that prevails over the density redistribution effect. This, in turn, leads to DC structures that yield similar bulk metallicities to CE, independently of the extent of the composition gradient.
  • When increasing the atmospheric metallicity in the DC structure to one (DCA) or three times stellar (DCA3), the inferred bulk metallicity becomes closer to the one yielded by FM. This demonstrates the importance of calibrating it with future observational data to better constrain internal metallicity estimates.
  • When considering non-adiabatic temperature gradients in DC structures, the inferred bulk metallicity can increase by up to 35% for a used value of $R_{\rho}=0.4$. However, future studies should explore a larger parameter space and assess the long-term stability of composition gradients.
  • Using a sample of 44 giant exoplanets, we confirm that heavy-element mass increases with planetary mass. The slope of the $M_{Z}-M$ relationship does not change with the assumed structure, but only the intercept.
@@
Future studies that account for additional physical processes such as convective mixing as well as the observed atmospheric compositions can improve giant planet characterisation. In addition, further investigations of breakpoints via multi-linear or orthogonal-distance regressions (Müller et al., 2024) in the inferred relations are desirable. Accurate measurements of the atmospheric chemistry of giant planets by JWST can also further constrain the planetary bulk composition and internal structure. In addition, forthcoming measurements from PLATO and Ariel (expected to be launched in 2026 and 2029, respectively) will sharpen bulk metallicity estimates by providing tighter constraints on mass, radius, age, and atmospheric composition for many planets. This will allow a better understanding of the interiors of giant planets and of their formation and evolution histories.
@@ Acknowledgements.
We thank Stefano Wirth, Simon Müller, Henrik Knierim, and Shay Zucker for insightful discussions.
Appendix A Planetary data
Table 5: List of retrieved adiabatic (DCA) and super-adiabatic runs for D structures (tested $R_{\rho}$alues of 0.1 and 0.4) for 11 planets.
Planet $R_{\rho}=0$ (DCA) $R_{\rho}=0.1$ $R_{\rho}=0.4$
HAT-P-12b 0.123 $\pm$ 0.047 0.1232 $\pm$ 0.043 0.189 $\pm$ 0.047
TOI-3629b 0.427 $\pm$ 0.029 0.4309 $\pm$ 0.0284 0.4752 $\pm$ 0.0249
HD332231b 0.210 $\pm$ 0.046 0.2222 $\pm$ 0.047 0.3127 $\pm$ 0.0651
HAT-P-19b 0.046 $\pm$ 0.017 0.0492 $\pm$ 0.022 0.055 $\pm$ 0.022
K2-287b 0.270 $\pm$ 0.023 0.2845 $\pm$ 0.0237 0.3306 $\pm$ 0.0252
HATS-49b 0.406 $\pm$ 0.032 0.4266 $\pm$ 0.0444 0.4342 $\pm$ 0.0194
HATS-75b 0.220 $\pm$ 0.023 0.2201 $\pm$ 0.0253 0.2735 $\pm$ 0.0287
HAT-P-17b 0.049 $\pm$ 0.026 0.050 $\pm$ 0.025 0.063 $\pm$ 0.036
WASP-84b 0.135 $\pm$ 0.045 0.1448 $\pm$ 0.0464 0.155 $\pm$ 0.045
HAT-P-54b 0.161 $\pm$ 0.049 0.1708 $\pm$ 0.0502 0.2054 $\pm$ 0.0529
TOI-1811b 0.141 $\pm$ 0.042 0.142 $\pm$ 0.042 0.177 $\pm$ 0.044
@@ Appendix A Planetary data
Figure 8: Output comparison between the data retrieved in this study, and the one found in Howard et al. (2025b). Upper panels: Bulk metallicity (normalised by the star’s metallicity) over mass for the initial 21 planets (as mentioned in Fig. LABEL:tab:sample_fullwidth_1 and LABEL:tab:sample_fullwidth_2) both for FM and CE structures. Bottom panels: Normalised difference (calculated with eq. 4) between the two estimates.
In order to assess the strength and accuracy of the retrieved $Z$ for the sample used, we took advantage of the fact that the used dataset is the same as the one used by Howard et al. (2025b). We thus took the initial sample of 21 exoplanets with the most precise data on Radius ($R_{err}$¡3%), and compared it with the same planets processed with the method explained in Howard et al. (2025b). Results are shown in Fig. 8, where a visual comparison between the trends for the $ZZ_{\star}$ versus $M$ is shown. To note how the best-fit lines vary from the ones retrieved with the full dataset of 44 exoplanets, showing the variability and unreliability of such curves for model prediction. Howard also discussed in their paper the shift of the slope $\alpha$ between the dataset, considering only $R_{err}$¡3%, or for the broader range considered of $R_{err}$¡5%. In the bottom panels of Fig. 8, it is shown how we assessed the difference between the data from this paper compared to the ones from Howard et al. (2025b). While a simple subtraction could provide information about the difference between the two mean values, we also wanted to investigate whether each ranges of values (so accounting for data uncertainty) lie within the other. To do so, a normalisation using the standard deviations of the data has been computed. The normalised difference can be expressed with the following formula:
$\sigma=\frac{\text{$CE/FM_{Howard}$}-\text{$CE/FM$}}{\sqrt{\sigma_{\text{Howard}}^{2}+\sigma_{\text{CE/FM}}^{2}}}.$ (4)
Results regarding this comparison were considered positive. While there is a slight change in the predictive trends between the two structures CE and FM (slight shift in $\beta$ for CE, and in $\alpha$ for FM compared to the ones found in Howard’s. The results for the normalised difference ultimately confirmed the similarities between the two retrieved results, and their respective methods; All values lie between a standard deviation of 1 and -1 (or $\sigma$—¡ 1), showing statistical consistency for all 21 cases compared. Furthermore, the normalised difference does not show any systematic error related to the mass (since there is no trend in the normalised difference $\sigma$). We thus conclude by saying that there is no discouraging difference between the two results (as the difference between the two approaches can be explained to be within the data error range and thus not due to randomness). We then accepted the validity of our approach, explained in Section 2.3, and proceeded with the inspection of the additional structures.
@@ Appendix C Model assumption-driven changes
Figure 9: Change in bulk metallicity according to the assumed structure and their difference relative to the FM model over mass.
@@
Appendix D WLS $\chi^{2}$ test
To assess the reliability of the power-law fits, we performed a WLS $\chi^{2}$ test, which accounts for individual data uncertainties rather than relying solely on $R^{2}$. In WLS, each residual is scaled by its measurement error: